OpenStudy (anonymous):

Show all work to solve 2 over x plus 1 over 9=1 over 3

OpenStudy (anonymous):

2/x + 1/9 = 1/3 (18+x)/9x = 1/3 3(18+x) = 9x 54+3x = 9x 54 = 9x - 3x 54 = 6x 9 = x